OpenStudy (anonymous):

in the first lecture ..the professor mentioned more notations..f'=(df/dx)=(dy/dx) and that deltaY =delta F,then he wrote (Omits X0) what does that mean?? omits x0?? and in the example of f(x)=x^n ,n=1,2,3,..... find d/dx (x^n) he said that when we substitute in the main formula for computing derivative we won't use x0 ,instead we will just write x as we have no X floating around different from the X0 afterwards he said that x is fixed and delta x is moving so shouldn't there be an x0 for the initial point and x for the final one?? and please someone explain the Binomial theorem :)

OpenStudy (anonymous):

I didn't get the first question properly , but here's what I know : here in my example I noted delta x with h, so when h->0 the value of the slope between the two points (the secant) goes to the value of the slope of the tangent we usually denote that by\[ \lim_{h \rightarrow 0} \frac{ \Delta y }{\Delta x }= \lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h }=\frac{ dy }{ dx }\]